OpenStudy (ememnim):

Find the coordinates of the midpoint of the line segment that joins the two given points. (x, y) = (−53, 19) and (x, y) = (19, −38)

OpenStudy (shadowlegendx):

There is a formula for this

OpenStudy (shadowlegendx):

Gimme a sec

OpenStudy (ememnim):

this is timed ): help

OpenStudy (shadowlegendx):

\[M = \frac{ (x _{1} + x _{2} ) }{ 2 } , \frac{ (y _{1} + y _{2} }{ 2 }\]

OpenStudy (shadowlegendx):

Input the coordinates, then solve

OpenStudy (shadowlegendx):

\[M = \frac{ -53 + 19 }{ 2 } , \frac{ 19 + -38 }{ 2 }\]

OpenStudy (shadowlegendx):

\[M = \frac{ -34 }{ 2 } , \frac{ -19 }{ 2 }\]

OpenStudy (shadowlegendx):

\[M = -17, -9.5\]

OpenStudy (shadowlegendx):

M = midpoint

OpenStudy (emilyhoot123):

should i put fraction or no

OpenStudy (shadowlegendx):

Is this multiple choice?

OpenStudy (emilyhoot123):

no fill in

OpenStudy (shadowlegendx):

I would gave it as -9.5, unless your teacher has instructed or showed otherwise

OpenStudy (shadowlegendx):

Reason is, -9.5 is easier to graph with, than -19/2

OpenStudy (shadowlegendx):

With -9.5 you just take the point in-between -9 and -10, with -19/2 you would go down 19 steps on the y axis from point (0,0), then cut that distance in half. tl;dr, -9.5 is the way to go

OpenStudy (shadowlegendx):

Looks good

OpenStudy (ememnim):

ok cool

OpenStudy (ememnim):